OpenStudy (anonymous):

need help will give medal +fan :-) Trigonometry. i have 5 questions what is the amplitude period and phase shift of f(x)=-3 cos (4x+pi)+6? a.) amplitude =3 , period = pi/2 , phase shift; x=-pi/4 b.) amplitude =-3, period =pi/2, phase shift : x=pi/4 c.)amplitude = -3 , period =-pi/2, phase shift : x=-pi /4 d.)amplitude =3 , period =2n, phase shift :x=-pi/4

OpenStudy (anonymous):

@TERM_THE_INATOR @ganeshie8 @rock_mit182 @secret66

OpenStudy (anonymous):

@beccaboo333 @thomaster @hartnn @HelpBlahBlahBlah

OpenStudy (anonymous):

@whpalmer4 @robtobey

OpenStudy (anonymous):

@RadEn @radar

OpenStudy (anonymous):

@crabby_patty

OpenStudy (anonymous):

i think it is b. am i right

OpenStudy (whpalmer4):

\[f(x) = A \cos(B(x + C)) +D\] Amplitude is \(A\) Period is \(\dfrac{2\pi}{B}\) Phase shift is \(C\) DC offset is \(D\)

OpenStudy (anonymous):

i know i just reread that in my book i justa wanted a second opinion do you think b is right? @whpalmer4

OpenStudy (whpalmer4):

so, looking at your equation: \[-3\cos(4x+\pi) + 6 = -3\cos (4(x+\frac{\pi}{4}) + 6\]\[-3\cos (4(x+\frac{\pi}{4}) + 6=A\cos(B(x+C))+D\]\[A = -3,B=4,C=\frac{\pi}{4},D=6\] so amplitude is \(-3\), period is \(\dfrac{2\pi}{4 }= \dfrac{\pi}{2}\), phase shift is \(\dfrac{\pi}{4}\) If that matches B, then B is correct.

OpenStudy (anonymous):

yes thank you its always nice to have a second opinion